OpenStudy (calculusxy):

The lines defined by Pt = (4+5t, −1+2t) and Qu = (4−2u, −1+5u) intersect perpendicularly. What are the coordinates of the point of intersection?

OpenStudy (calculusxy):

@agent0smith

OpenStudy (agent0smith):

You could convert from the parametric equations back to a linear equation, then just find where the two lines intersect.

OpenStudy (calculusxy):

could i do y = 2/5x -1 (Pt) y = -5/2x - 1

OpenStudy (calculusxy):

y = -5/2x - 1 (Qu)

OpenStudy (calculusxy):

are the equations correct?

OpenStudy (agent0smith):

Sorry, gotta go

OpenStudy (calculusxy):

ok thanks for your help.

OpenStudy (calculusxy):

@zepdrix can you help me?

zepdrix (zepdrix):

I'm a little confused. Are these vectors? \(\large\rm \vec P(t) = <4+5t, −1+2t>\) Or just coordinates using some parameter? \(\large\rm P(t)=(4+5t,-1+2t)\) Ahh my brain >.<

OpenStudy (calculusxy):

\[P_t = (4+5t, -1+2t)\] I think they are parameters.

OpenStudy (eliesaab):

Notice that P[0]=Q[0]=(4,-1) That is the point of intersectio

OpenStudy (calculusxy):

Oh Wow! Thank you @eliesaab

OpenStudy (eliesaab):

YW @calculusxy

OpenStudy (calculusxy):

@eliesaab Would the equation for line Pt y = 2/5x - 1 be incorrect?

OpenStudy (eliesaab):

No, The right equation is \[y=\frac{1}{5} (2 x-13)\]